OpenStudy (jasmines8):

How do you make equations for the following problem: A cell phone provider offers a plan that costs $40 per month plus $.20 per text message sent or received. A compared plan costs $60 per month but offers unlimited text messaging.

OpenStudy (mathstudent55):

Let's use y to mean the monthly cost. For the second plan, what is the monthly cost (using y in an equation)?

OpenStudy (jasmines8):

$60x

OpenStudy (mathstudent55):

What is x?

OpenStudy (mathstudent55):

Also, where is y?

OpenStudy (jasmines8):

Thts the problem I'm having idk if x should b monthly cost or texts

OpenStudy (mathstudent55):

Here is the second plan: "A compared plan costs $60 per month but offers unlimited text messaging." The cost of the plan is simply $60 per month. Since it has unlimited texts, it makes no difference how many texts you use in a month. Al texts are included.

OpenStudy (mathstudent55):

I started with the second plan because it is easier to write an equation for since it is a simpler plan.

OpenStudy (jasmines8):

Yes but idk how to add $.20 per text if x is already for monthly cost

OpenStudy (jasmines8):

Ok

OpenStudy (mathstudent55):

Let y = monthly cost. Let x = number of texts per month. Second plan: y = 60

OpenStudy (jasmines8):

Ok but wouldn't it b y=60x?

OpenStudy (mathstudent55):

Do you understand why the second plan is simply y = 60? It is because no matter how many texts are used, the price is always $60 per month.

OpenStudy (jasmines8):

Yea I do understand Tht part

OpenStudy (mathstudent55):

x is the number of texts used in one month. The second plan has unlimited texts, so the price of each text is $0. You could write the equation for the second text as y = 60 + 0x which means the cost is $60 plus $0 per text. Since $0 per text will end up being $0 no matter how many texts are sent, you simplify y = 60 + 0x to just y = 60

OpenStudy (jasmines8):

Oh! So the equation for the first one would b y=40+.25x??

OpenStudy (mathstudent55):

Now let's look at the first plan. The first plan also has a fixed monthly amount. It is $40 instead of $60. Now you do need to pay for each text, and each text costs $0.20. If you use x number texts, the x number of texts will cost you x * 0.20 which can be written as 0.20x. That means the firs plan costs: y = 40 + 0.20x

OpenStudy (mathstudent55):

You're on the right track, but the texts cost $0.20 each, not $0.25 each, using the first plan..

OpenStudy (jasmines8):

yea Thts wht I meant Idk y I added the 5

OpenStudy (mathstudent55):

After all the discussion, here is the answer: First plan: y = 40 + 0.2x Second plan: y = 60

OpenStudy (mathstudent55):

I think you got it now.

OpenStudy (jasmines8):

Yes thanks do u think u could help me with one more probables like Tht one it's the last one I need to find

OpenStudy (mathstudent55):

Sure. Pls start a new post.

OpenStudy (jasmines8):

ok